# A Survey of Low-bit Large Language Models: Basics, Systems, and Algorithms

### v2

@@ Abstract
Large Language Model, Quantization, Low-bit, System, Algorithm
journal: Neural Networks
@@
[beihang]organization=Beihang University, addressline=37 Xueyuan Road, city=Haidian District, postcode=100191, state=Beijing, country=China
[…]
[eth]organization=ETH Zurich, addressline=Rämistrasse 101, city=Zurich, postcode=8092, state=Zurich, country=Switzerland
[…]
[sensetime]organization=SenseTime, addressline=No.1900 Hongmei Road, city=Xuhui District, postcode=200233, state=Shanghai, country=China
[…]
[cuhk]organization=The Chinese University of Hong Kong, addressline=Shatin, city=NT, postcode=999077, state=Hong Kong SAR, country=China
@@ 1 Introduction
Large language models (LLMs) (OpenAI et al., 2024; Touvron et al., 2023a, b; Dubey et al., 2024; Lozhkov et al., 2024; Liu et al., 2024a) have revolutionized natural language processing by delivering unprecedented performance across a range of tasks, from text generation to language understanding. However, their remarkable capabilities come with significant computational and memory demands. This has raised considerable challenges when deploying these models in scenarios with limited resources or high concurrency. To address these challenges, low-bit quantization has emerged as a pivotal approach for enhancing the efficiency and deployability of LLMs.
@@ 2.1.1 Standard Formats
Floating-point Numbers. The floating-point data type is comprehensively defined in the IEEE 754 (IEEE, 2019) standard, which is also the most prevailing number format in computer systems. Let us denote them as $\mathrm{FP}k$, where $k$ represents the number of bits that the value occupies in memory, usually $32,16,8$, etc. A floating-point number can be uniformly expressed as:
$X_{\mathrm{FP}k}=(-1)^{s}2^{p-bias}(1.\texttt{mantissa})=(-1)^{s}2^{p-bias}(1+\frac{d_{1}}{2}+\frac{d_{2}}{2^{2}}+...+\frac{d_{m}}{2^{m}}),$ (1)
@@
where $s$ is the sign bit, $p$ is the exponent integer, $bias$ is applied to the exponent, $m$ is the total number of mantissa bits in the significand, and $d_{1},d_{2},\ldots,d_{m}$ represent the digits of the mantissa part in the binary format. The bits of $s$, $p$ and $m$ should be accumulated to $k$ for an $\mathrm{FP}k$ value.
[…]
@@
NormalFloat (NF) (Dettmers et al., 2024) is a fixed floating-point method used in weight-only quantization strategies for LLMs. The data representing format follows the floating-points, but the $2^{k}$ values $X^{\mathrm{NF}}_{i},i\in[0,2^{k}-1]$ are estimated to be:
$X^{\mathrm{NF}}_{i}=\frac{1}{2}\left(quantile\left(N(0,1),\frac{i}{2^{k}+1}\right)+quantile\left(N(0,1),\frac{i+1}{2^{k}+1}\right)\right),$ (2)
@@
where $quantile(\cdot,q)$ is the $q$-th quantiles of the input. $N(0,1)$ means the standard normal distribution. For a tensor that does not fall within the range of -1 to 1, we must first scale it using its maximum absolute value. To ensure the exact representation for zero, it asymmetrically divides the data into the positive and negative parts by estimating $2^{k-1}$ of $X^{\mathrm{NF}}_{i}$ for the negative and $12^{k-1}-$ 1 for the positive, then removes one of the zeros in both sets. NF is estimated to have an almost equal expected number of values in each quantization bin to keep the most information in the quantized formats.
@@
Micro Scaling FP (Rouhani et al., 2023). It was proposed and developed in collaboration with industry alliance members, including AMD, Arm, Intel, Meta, Microsoft, NVIDIA, and Qualcomm, which aims to establish a unified standard for fine-grained sub-blocks of tensor format. It applies E8M0 scaling factors on a block of data with various original formats (i.e., FP8, FP6, FP4, INT8). The scaling block size indicates the number of elements that each scaling applies. It keeps high precision for the value representation but is significantly efficient on hardware due to the shared scalings.
@@
where $m=k-1$ and $s\in\{0,1\}$ for signed integers. $m=k$ for unsigned integers while we regard $s=0$. Therefore, the signed integers range from $[-2^{k-1},2^{k-1}-1]$, and the unsigned one $[0,2^{k}-1]$.
Binarized Numbers. Binarization is the most aggressive quantization technique, which directly abstracts the sign of value (Liu et al., 2018; Qin et al., 2022; Li et al., 2024d). It will lose most information, but bring significant acceleration and parameter compression in inference. The hardware takes $0,1$ for each bit originally, but developers define different logic rules and accumulation algorithms to achieve various binarized computations. Therefore, floating-point numbers can be binarized to $\{-1,1\}$ or $\{0,1\}$, depending on what value we expect the single bit to represent in our algorithms.
@@
Table 1: Min and Max values for different number formats (IEEE, 2019).
@@ 2.1.2 Custom Formats
For faster computation and better fitting the numerical distributions of LLMs, many studies propose custom number formats besides the standard formats described above. Here we introduce three typical customized formats. We omit the works before the advent of LLMs (Tambe et al., 2019) because their performance has not been validated on LLMs.
Floating-point Integer (Flint) (Guo et al., 2022) combines the advantages of floating-point and integer representations, which is $X_{\mathrm{Flint}}=2^{{p}-bias}\times(1.\texttt{mantissa})$. We take the 4-bit Flint on float-based MAC units as an example:
$p=\left\{\begin{array}[]{ll}3-\texttt{LZD}(b_{2}b_{1}b_{0}),&\quad b_{3}=0\\ 4+\texttt{LZD}(b_{2}b_{1}b_{0}),&\quad b_{3}=1\end{array},\right.\quad\texttt{mantissa}=b_{2}b_{1}b_{0}\texttt{<<}(\texttt{LZD}(b_{2}b_{1}b_{0})+1),$ (4)
where the LZD denotes the Leading Zero Detector (Oklobdzija, 1994) which accumulates the leading zeros on the left of the bitstring, << is the left shift operation, and $bias=1$ for float-based Flint4. It expands the range by integrating exponents into the integers, therefore. Compared to pure integers, Flint can represent a larger range with a limited number of bits, which better fits the distribution of LLM parameters.
Adaptive Biased Float (Abfloat) is first proposed in Outlier-Victim Pair Quantization (OVP) (Guo et al., 2023a) to deal with outliers. The difference to Flint is that Abfloat applies a bigger $bias$ to the exponent, and left shifts $m$-bit to enlarge the $1$ before mantissa, making the magnitude even larger to cover the outliers. The $\mathrm{E}e\mathrm{M}m$ Abfloat value can be expressed as:
@@
[…]
Student Float (SF) (Dotzel et al., 2024a) follows the floating-point format but has specific fixed points for quantization, which is different from the above two types. SF is an improvement of NF in Section 2.1.1 and holds the view that the parameters obey Student’s t-distribution $S(t;\nu)$, of which the probability density function is:
@@
where $\omega=\frac{1}{2}(\frac{1}{32}+\frac{1}{30})$, $\{q_{1},\dots,q_{8}\}$ and $\{q_{9},\dots,q_{16}\}$ are two groups of evenly spaced quantiles. Then we normalize $\tilde{X}^{\mathrm{SF}}$ to $[-1,1]$ by ${X}^{\mathrm{SF}}_{i}=\frac{\tilde{X}^{\mathrm{SF}}_{i}}{\max_{i}|\tilde{X}^{\mathrm{SF}}_{i}|}$. As $\nu$ increases, the peaks of the t-distribution become shorter and wider, and SF4 spreads out more. It converges to the standard normal distribution (NF) as $\nu\to\infty$. Same as NF, SF is used in weight-only quantization (which we introduce in Section 3.2.1). Therefore, it does not need the low-level definition of base operations but requires a custom dequantization from SF to standard formats.
@@ 2.2 Quantization Granularity
Tensor-wise is the simplest and coarsest granularity, which takes a single scaling factor and zero point to the entire tensor (Zhang et al., 2024c). It can be the fastest but may lead to the most performance degradation because it is incapable of handling the values with a wide variation. Therefore, it is unsuitable for cases where accuracy is important or the task/model is sensitive to quantization.
Token-wise is used in LLMs only, which means that each token (word or subword) has a scaling (Yao et al., 2022). It captures the fine-grained variations in different tokens. Usually, we adopt dynamic token-wise quantization for activation to reduce the quantization error and ensure diversity in generative models.
Channel-wise means each channel in weight within a tensor uses one scale and can be merged into quantized weight (Kim et al., 2024). Token-wise activation and channel-wise weight are usually used together. Because for $i$-th token in activation and $j$-th channel in weight, the corresponding $s_{\textbf{x}_{i}}\in s_{\textbf{x}}\in\mathbb{R}^{T\times 1}$ and $s_{\textbf{w}_{j}}\in s_{\textbf{w}}\in\mathbb{R}^{1\times C}$ can be calculated first as $s\in\mathbb{R}^{[1]}$ and multiplied to the coordinate $[i,j]$ in output matrix $\textbf{X}_{O}$. In this way, we preserve the generation performance with little computation overheads.
Group-wise balances the computational complexity and the quantization error by grouping tensors or channels with the same scaling factor. It also reduces the storage of scaling factors by $g$ if there are $g$ tokens/channels per group (Heo et al., 2023; Yao et al., 2022).
Element-wise is only applied while training the weight, which is always used together with another quantization granularity, such as tensor-wise (see Figure 2(e)). Before inference, the element-wise scaling is merged into the quantized weight. Therefore, only the tensor-wise scale is needed to be computed in inference (Lee et al., 2023) to recover the value magnitude.
Different quantization granularity are always combined and adopted together. For example, Lee et al. (2023) uses a channel-wise scale for the Key matrix but a token-wise scale for the Value matrix based on the distribution of the data. More algorithms can be found in Section 5.2.3.
@@ 2.3 Dynamic and Static Quantization
Dynamic Quantization (Krishnamoorthi, 2018; Liu et al., 2022) calibrates and stores quantized weight. Usually, it does not need input data, but searches for the optimal scaling factors $s_{\mathbf{w}}$ and zero-points $Z_{\mathbf{w}}$ by minimizing the quantization error for each tensor of weight. During inference, the activation will be input into the quantization module to compute the optimal scaling factors $s_{\mathbf{x}}$ and $Z_{\mathbf{x}}$, and then quantized to INT8 by the dynamically computed factors before conducting integer GEMM with quantized weight. The scaling and zero point of activation are obtained in real time based on the current batch of input data. Therefore, the scaling factor flexibly adapts the input data distribution, bringing the smallest quantization error. While it takes extra computational complexity to get the scale during inference. It is suitable for scenarios that require rapid deployment because it does not require calibration.
Static Quantization (Bai et al., 2021) takes calibration data consisting of a small fraction of the training dataset. By inputting the samples into the model, we find the optimal scaling factors for both weight and activation (the middle one in Figure 3) or weight only (the right one), and are fixed during inference. It allows for the evaluation of the quantized model during preparation, ensuring that quantization does not significantly harm the model’s performance. As for inference, the middle one in Figure 3 quantizes the activation to low-bit and computes low-bit GEMM (Dettmers et al., 2022b) with quantized weight. For the right one in Figure 3, the weight will be dequantized to floating-point numbers, and the activation will not be quantized before conducting floating-point GEMM (Lin et al., 2024a), thus we name it weight-only quantization.
@@ 3.1 Inference Framework for Quantization
We list the representative inference frameworks in Table 2. Currently, no single inference framework dominates in terms of performance or usage. However, some classic deep learning frameworks, such as TensorRT-LLM [^1] https://github.com/NVIDIA/TensorRT-LLM, ONNX-runtime [^2] https://github.com/microsoft/onnxruntime, Transformers [^3] https://huggingface.co/docs/transformers/en/index (Huggingface), OpenVINO [^4] https://github.com/openvinotoolkit/nncf, PowerInfer [^5] https://github.com/SJTU-IPADS/PowerInfer, PPLNN [^6] https://github.com/openppl-public/ppl.nn, and Xorbits Inference [^7] https://github.com/xorbitsai/inference have integrated the support for efficient inference of large models. In addition, other inference frameworks emerged after the advent of large models that are specifically proposed for LLMs, such as bitsandbytes [^8] https://github.com/bitsandbytes-foundation/bitsandbytes, ctransformers [^9] https://github.com/marella/ctransformers, MLC-LLM [^10] https://github.com/mlc-ai/mlc-llm, DeepSpeed-MII [^11] https://github.com/microsoft/DeepSpeed-MII, vLLM [^12] https://github.com/vllm-project/vllm, LMDeploy [^13] https://github.com/InternLM/lmdeploy, LightLLM [^14] https://github.com/ModelTC/lightllm, QServe [^15] https://github.com/mit-han-lab/qserve, llama.cpp [^16] https://github.com/ggerganov/llama.cpp, llama2.c [^17] https://github.com/karpathy/llama2.c, inferflow [^18] https://github.com/inferflow/inferflow, ScaleLLM [^19] https://github.com/vectorch-ai/ScaleLLM, SGLang [^20] https://github.com/sgl-project/sglang, gpt-fast [^21] https://github.com/pytorch-labs/gpt-fast, FastChat [^22] https://github.com/lm-sys/FastChat, OpenLLM [^23] https://github.com/bentoml/OpenLLM and so on. These frameworks are lightweight and have integrated many specialized optimization techniques for large models.

### v3

@@ 1 Introduction
Large language models (LLMs) (openai2024gpt4; touvron2023llama; touvron2023llama2; dubey2024llama; lozhkov2024starcoder; liu2024deepseek) have revolutionized natural language processing by delivering unprecedented performance across a range of tasks, from text generation to language understanding. However, their remarkable capabilities come with significant computational and memory demands. This has raised considerable challenges when deploying these models in scenarios with limited resources or high concurrency. To address these challenges, low-bit quantization has emerged as a pivotal approach for enhancing the efficiency and deployability of LLMs.
@@ 2.1.1 Standard Formats
Floating-point Numbers. The floating-point data type is comprehensively defined in the IEEE 754 (ieee754) standard, which is also the most prevailing number format in computer systems. Let us denote them as $\mathrm{FP}k$, where $k$ represents the number of bits that the value occupies in memory, usually $32,16,8$, etc. A floating-point number can be uniformly expressed as:
$X_{\mathrm{FP}k}$ $=$ ${(-1)}^{s}2^{p-bias}(1.\texttt{mantissa})={(-1)}^{s}2^{p-bias}$ (1)
$\left(1+\frac{d_{1}}{2}+\frac{d_{2}}{2^{2}}+\ldots+\frac{d_{m}}{2^{m}}\right),$
@@
where $s$ is the sign bit, $p$ is the exponent integer, biasbias is applied to the exponent, $m$ is the total number of mantissa bits in the significand, and $d_{1},d_{2},\ldots,d_{m}$ represent the digits of the mantissa part in the binary format. The bits of $s$, $p$ and $m$ should be accumulated to $k$ for an $\mathrm{FP}k$ value.
Since LLMs occupy more memory, lower-bit formats become popularly adopted in both training and inference. We omit the 32-bit number format here since the 16 and lower bitwidth has become the mainstream practice in application. We can further categorize each $\mathrm{FP}k$ according to its bit allocations for the exponent (E) and mantissa (M) parts. We use $\mathrm{E}e\mathrm{M}m$ to denote the subcategories. As for $\mathrm{FP}16$, IEEE 754 defines float16 (also known as half-precision or FP16) and bfloat16 (brain floating point or BF16), which can be represented as E5M10 and E8M7, respectively. Therefore, bfloat16 can represent larger magnitudes with more exponent bits (identical to that of FP32) while more sparse than float16 with less mantissa in the significand, which may exert unprecedented potential in LLMs (henry2019leveragingbfloat16artificialintelligence). As well as E4M3 and E5M2 for $\mathrm{FP}8$, both are standard formats that are already supported by several mainstream deep learning inference engines, such as MLC-LLM, Quanto, and so on (see Section 3.1.2 for details).
@@
NormalFloat (NF) (dettmers2024qlora) is a fixed floating-point method used in weight-only quantization strategies for LLMs. The data representing format follows the floating-points, but the $2^{k}$ values $X^{\mathrm{NF}}_{i},i\in[0,2^{k}-1]$ are estimated to be:
$X^{\mathrm{NF}}_{i}$ $=\frac{1}{2}\Bigl(quantile\!\left(N(0,1),\frac{i}{2^{k}+1}\right)$
$\quad+\,quantile\!\left(N(0,1),\frac{i+1}{2^{k}+1}\right)\Bigr),$ (2)
@@
where $quantile(\cdot,q)$ is the $q$-th quantiles of the input. $N(0,1)$ means the standard normal distribution. For a tensor that does not fall within the range of -1 to 1, we must first scale it using its maximum absolute value. To ensure the exact representation for zero, it asymmetrically divides the data into the positive and negative parts by estimating $2^{k-1}$ of $X^{\mathrm{NF}}_{i}$ for the negative and $2^{k-1}-1$ for the positive, then removes one of the zeros in both sets. NF is estimated to have an almost equal expected number of values in each quantization bin to keep the most information in the quantized formats.
@@
Micro Scaling FP (rouhani2023microscalingdataformatsdeep). It was proposed and developed in collaboration with industry alliance members, including AMD, Arm, Intel, Meta, Microsoft, NVIDIA, and Qualcomm, which aims to establish a unified standard for fine-grained sub-blocks of tensor format. It applies E8M0 scaling factors on a block of data with various original formats (i.e., FP8, FP6, FP4, INT8). The scaling block size indicates the number of elements that each scaling applies. It keeps high precision for the value representation but is significantly efficient on hardware due to the shared scalings.
@@
where m=k1m=k-1 and $s\in\{0,1\}$ for signed integers. $m=k$ for unsigned integers while we regard $s=0$. Therefore, the signed integers range from $2^{k-1},2^{k-1}-1]$, and the unsigned one $0,2^{k}-1]$. Before the advent of LLMs, integer quantization had been applied in BERT-based language models, as demonstrated by qbert.
Binarized Numbers. Binarization is the most aggressive quantization technique, which directly abstracts the sign of value (birealnet; bibert; bi-vit). It will lose most information, but bring significant acceleration and parameter compression in inference. The hardware takes $0,1$ for each bit originally, but developers define different logic rules and accumulation algorithms to achieve various binarized computations. Therefore, floating-point numbers can be binarized to $\{-1,1\}$ or $\{0,1\}$, depending on what value we expect the single bit to represent in our algorithms. Some studies further extended binarization to ternary quantization. Before the emergence of LLMs, works such as bai-etal-2021-binarybert; zhang-etal-2020-ternarybert; bit; liu-etal-2023-binary explored binary or ternary quantization formats.
@@
Table 1: Min and Max values for different number formats (ieee754).
@@ 2.1.2 Custom Formats
For faster computation and better fitting the numerical distributions of LLMs, many studies propose custom number formats besides the standard formats described above. Here we introduce three typical customized formats. We omit the works before the advent of LLMs (adaptivefloat) because their performance has not been validated on LLMs.
Floating-point Integer (Flint) (Flint) combines the advantages of floating-point and integer representations, which is $X_{\mathrm{Flint}}=2^{{p}-bias}\times(1.\texttt{mantissa})$. We take the 4-bit Flint on float-based MAC units as an example:
$p=$ $\left\{\begin{array}[]{@{}l@{}l@{}l@{}}3-\texttt{LZD}(b_{2}b_{1}b_{0}),&\qquad b_{3}=0\\ 4+\texttt{LZD}(b_{2}b_{1}b_{0}),&\qquad b_{3}=1\end{array},\right.\qquad\texttt{mantissa}$ (5)
$=b_{2}b_{1}b_{0}\texttt{<<}(\texttt{LZD}(b_{2}b_{1}b_{0})+1),$ (6)
where the LZD denotes the Leading Zero Detector (oklobdzija1994algorithmic) which accumulates the leading zeros on the left of the bitstring, << is the left shift operation, and $bias=1$ for float-based Flint4. It expands the range by integrating exponents into the integers, therefore. Compared to pure integers, Flint can represent a larger range with a limited number of bits, which better fits the distribution of LLM parameters.
Adaptive Biased Float (Abfloat) is first proposed in Outlier-Victim Pair Quantization (OVP) (Abfloat) to deal with outliers. The difference to Flint is that Abfloat applies a bigger biasbias to the exponent, and left shifts $m$-bit to enlarge the $1$ before mantissa, making the magnitude even larger to cover the outliers. The $\mathrm{E}e\mathrm{M}m$ Abfloat value can be expressed as:
@@
When $\text{bias}=0$, the range is similar to $\text{Flint}4$. With $\text{bias}=2$ for E2M1, the range changes to $\{12,\dots,96\}$. With $\text{bias}=3$, the range further extends to $\{24,\dots,192\}$. The other difference to Flint is that Abfloat is only adopted on outliers, but the normal values are stored in INT4/8 or Flint4. Both data formats require custom system support to define the behavior of the base operations (such as addition, multiplication, and so on).
Student Float (SF) (studentFormat) follows the floating-point format but has specific fixed points for quantization, which is different from the above two types. SF is an improvement of NF in Section 2.1.1 and holds the view that the parameters obey Student’s t-distribution $S(t;\nu)$, of which the probability density function is:
@@
where $\omega=\frac{1}{2}(\frac{1}{32}+\frac{1}{30})$, $\{q_{1},\dots,q_{8}\}$ and $\{q_{9},\dots,q_{16}\}$ are two groups of evenly spaced quantiles. Then we normalize $\tilde{X}^{\mathrm{SF}}$ to [1,1][-1,1] by ${X}^{\mathrm{SF}}_{i}=\frac{\tilde{X}^{\mathrm{SF}}_{i}}{\max_{i}|\tilde{X}^{\mathrm{SF}}_{i}|}$. As $\nu$ increases, the peaks of the t-distribution become shorter and wider, and SF4 spreads out more. It converges to the standard normal distribution (NF) as $\nu\to\infty$. Same as NF, SF is used in weight-only quantization (which we introduce in Section 3.2.1). Therefore, it does not need the low-level definition of base operations but requires a custom dequantization from SF to standard formats.
@@ 2.2 Quantization Granularity
Tensor-wise is the simplest and coarsest granularity, which takes a single scaling factor and zero point to the entire tensor (zhang2024flattenquantbreakinginferencecomputebound). It can be the fastest but may lead to the most performance degradation because it is incapable of handling the values with a wide variation. Therefore, it is unsuitable for cases where accuracy is important or the task/model is sensitive to quantization.
Token-wise is used in LLMs only, which means that each token (word or subword) has a scaling (yao2022zeroquantefficientaffordableposttraining). It captures the fine-grained variations in different tokens. Usually, we adopt dynamic token-wise quantization for activation to reduce the quantization error and ensure diversity in generative models.
Channel-wise means each channel in weight within a tensor uses one scale and can be merged into quantized weight (kim2024memory). Token-wise activation and channel-wise weight are usually used together. Because for $i$-th token in activation and $j$-th channel in weight, the corresponding $s_{\textbf{x}_{i}}\in s_{\textbf{x}}\in\mathbb{R}^{T\times 1}$ and $s_{\textbf{w}_{j}}\in s_{\textbf{w}}\in\mathbb{R}^{1\times C}$ can be calculated first as $s\in\mathbb{R}^{[1]}$ and multiplied to the coordinate $[i,j]$ in output matrix $\textbf{X}_{O}$. In this way, we preserve the generation performance with little computation overheads.
Group-wise balances the computational complexity and the quantization error by grouping tensors or channels with the same scaling factor. It also reduces the storage of scaling factors by $g$ if there are $g$ tokens/channels per group (heo2023rethinking; yao2022zeroquantefficientaffordableposttraining).
Element-wise is only applied while training the weight, which is always used together with another quantization granularity, such as tensor-wise (see Figure 2(e)). Before inference, the element-wise scaling is merged into the quantized weight. Therefore, only the tensor-wise scale needs to be computed in inference (lee2023flexround) to recover the value magnitude.
Different quantization granularity are always combined and adopted together. For example, lee2023flexround uses a channel-wise scale for the Key matrix but a token-wise scale for the Value matrix based on the distribution of the data. More algorithms can be found in Section 5.2.3.
@@ 2.3 Dynamic and Static Quantization
Dynamic Quantization (krishnamoorthi2018quantizingdeepconvolutionalnetworks; Liu_2022_CVPR) calibrates and stores quantized weight. Usually, it does not need input data, but searches for the optimal scaling factors $s_{\mathbf{w}}$ and zero-points $Z_{\mathbf{w}}$ by minimizing the quantization error for each tensor of weight. During inference, the activation will be input into the quantization module to compute the optimal scaling factors $s_{\mathbf{x}}$ and $Z_{\mathbf{x}}$, and then quantized to INT8 by the dynamically computed factors before conducting integer GEMM with quantized weight. The scaling and zero point of activation are obtained in real time based on the current batch of input data. Therefore, the scaling factor flexibly adapts the input data distribution, bringing the smallest quantization error. While it takes extra computational complexity to get the scale during inference. It is suitable for scenarios that require rapid deployment because it does not require calibration.
Static Quantization (bai2021efficientposttrainingquantizationpretrained) takes calibration data consisting of a small fraction of the training dataset. By inputting the samples into the model, we find the optimal scaling factors for both weight and activation (the middle one in Figure 3) or weight only (the right one), and are fixed during inference. It allows for the evaluation of the quantized model during preparation, ensuring that quantization does not significantly harm the model’s performance. As for inference, the middle one in Figure 3 quantizes the activation to low-bit and computes low-bit GEMM (dettmers2022llmint8) with quantized weight. For the right one in Figure 3, the weight will be dequantized to floating-point numbers, and the activation will not be quantized before conducting floating-point GEMM (lin2024awq), thus we name it weight-only quantization.
@@ 3.1 Inference Framework for Quantization
We list the representative inference frameworks in Table 2. The inference process of Large Language Models (LLMs) consists of two key stages: Prefill and Decode. During the Prefill stage, the input prompt is tokenized and processed through the model’s Transformer layers to generate contextual embeddings, leveraging self-attention mechanisms to capture dependencies between tokens. This stage establishes a rich contextual representation of the input, which is stored for subsequent text generation. In the Decode stage, the model generates text autoregressively, predicting one token at a time by iteratively considering the sequence of previously generated tokens. This involves embedding lookup, attention computation, and token selection based on probability distributions. While the prefill stage processes the entire input at once, making it computationally intensive, the decode stage operates incrementally, building the output sequentially. Together, these stages enable LLMs to produce coherent and contextually relevant text, forming the foundation for optimization techniques like quantization, which aim to enhance efficiency without compromising performance. Currently, no single inference framework dominates in terms of performance or usage. However, some classic deep learning frameworks, such as TensorRT-LLM [^1] https://github.com/NVIDIA/TensorRT-LLM, ONNX-runtime [^2] https://github.com/microsoft/onnxruntime, Transformers [^3] https://huggingface.co/docs/transformers/en/index (Huggingface), OpenVINO [^4] https://github.com/openvinotoolkit/nncf, PowerInfer [^5] https://github.com/SJTU-IPADS/PowerInfer, PPLNN [^6] https://github.com/openppl-public/ppl.nn, and Xorbits Inference [^7] https://github.com/xorbitsai/inference have integrated the support for efficient inference of large models. In addition, other inference frameworks emerged after the advent of large models that are specifically proposed for LLMs, such as bitsandbytes [^8] https://github.com/bitsandbytes-foundation/bitsandbytes, ctransformers [^9] https://github.com/marella/ctransformers, MLC-LLM [^10] https://github.com/mlc-ai/mlc-llm, DeepSpeed-MII [^11] https://github.com/microsoft/DeepSpeed-MII, vLLM [^12] https://github.com/vllm-project/vllm, LMDeploy [^13] https://github.com/InternLM/lmdeploy, LightLLM [^14] https://github.com/ModelTC/lightllm, QServe [^15] https://github.com/mit-han-lab/qserve, llama.cpp [^16] https://github.com/ggerganov/llama.cpp, llama2.c [^17] https://github.com/karpathy/llama2.c, inferflow [^18] https://github.com/inferflow/inferflow, ScaleLLM [^19] https://github.com/vectorch-ai/ScaleLLM, SGLang [^20] https://github.com/sgl-project/sglang, gpt-fast [^21] https://github.com/pytorch-labs/gpt-fast, FastChat [^22] https://github.com/lm-sys/FastChat, OpenLLM [^23] https://github.com/bentoml/OpenLLM and so on. These frameworks are lightweight and have integrated many specialized optimization techniques for large models.